# MMAPG: A Training-Free Framework for Multimodal Multi-hop Question Answering via Adaptive Planning Graphs

### v1

@@ Abstract
Yiheng Hu1, Xiaoyang Wang1, Qing Liu 2, Xiwei Xu 2, Qian Fu 2, Wenjie Zhang1, Liming Zhu 2, 1University of New South Wales, Sydney, Australia 2CSIRO Data61, Australia {yiheng.hu, xiaoyang.wang1, wenjie.zhang}@unsw.edu.au {Q.Liu, Xiwei.Xu, Qian.Fu, Liming.Zhu}@data61.csiro.au
@@ 1 Introduction
[…]
@@
[…]
@@ 3.4 Retrieval Module
[…]
@@ 4.2 Main Results
To assess our method in long-range reasoning, we compare its performance with an iterative framework, SKURG in Figure 4, which shows the F1 score across varying numbers of required supporting contexts in MultimodalQA.
As the number of supporting contexts increases, requiring more steps, SKURG’s performance drops significantly, especially beyond five contexts. In contrast, our model maintains stable performance, which shows its robustness in complex reasoning scenarios.
@@ 4.3 Ablation Study
[…]
@@ 4.4 Case Study
We show a case study in Figure 5, which presents how the Adaptive Planning Graph is constructed. At the initial few steps, appropriate nodes are generated. For the third retrieval node, $R_{3}$, the system attempts to retrieve information regarding common profession, but mistakenly returns results related to Britney Spears without identifying her profession, leading to a deviation from the intended path. As Britney Spears’ profession is identified as an essential information in the overall plan, the planning module analyzes the graph state at this point and determines that the last retrieval action does not capture this information. Consequently, it decides to focus on retrieving Britney Spears’ profession directly in the next step. Ultimately, this leads to the correct answer. The case study highlights that, even when the graph temporarily deviates, the system is capable of recovering and producing accurate results.
@@ Appendix A Appendices
Figure 6: Comparison of similarity score distributions.
Figure 7: Projection of image and text embeddings.
A.1 Challenges of Modality-Agnostic Retrieval
[…]
A.2 Overall Workflow
A.3 Detailed Prompts
[…]
@@ A.5 Implementation
[…]

### v2

@@ Abstract
Yiheng Hu1,2, Xiaoyang Wang1, Qing Liu 2, Xiwei Xu 2, Qian Fu 2, Wenjie Zhang1, Liming Zhu 2, 1University of New South Wales, Sydney, Australia 2CSIRO Data61, Australia {yiheng.hu, xiaoyang.wang1, wenjie.zhang}@unsw.edu.au {Q.Liu, Xiwei.Xu, Qian.Fu, Liming.Zhu}@data61.csiro.au
@@ 1 Introduction
The second mainstream method employs an iterative approach Trivedi et al. (2022); Yang et al. (2023a); Zhang et al. (2024), which offers greater flexibility as it has no restrictions on the number of steps, allowing for the integration of additional sources as needed. The structure of these methods follow a single-path paradigm, where actions proceed in a fixed pattern. This design introduces cascading error propagation, which relies on the model’s inherent reasoning ability to correctify. For instance, in Figure 2, if an initial retrieval incorrectly identifies details about the TV show Pokémon, this mistake could mislead the next subquestion, ultimately leading to an incorrect result along a single-path with flaw. The challenge is further amplified in multimodal settings, where the modality involved in each step is often unknown. To handle different types of sources, SKURG Yang et al. (2023a) fuses image and text embeddings to create an entity-centered representation, which relies on extensive training of existing models. ETG Zhang et al. (2024) retrieves evidences by converting images into texts during preprocessing, which may result in information loss. and omit relevant contexts. These limitations highlight the need for developing methods that can dynamically adapt to multimodal sources.
@@
    We introduce a novel framework, MMAPG (Multimodal Multi-hop Adaptive Planning Graph), that offers enhanced flexibility for tackling multimodal multi-hop QA. By constructing the Adaptive Planning Graph step-by-step, our approach facilitates dynamic exploration of different sources and supports a graph-based reasoning flow. To the best of our knowledge, this is the first work using graph-based planning for multimodal multi-hop QA.
@@ 2.2 Chain of thought
Despite these advances, these reasoning methods might overlook the integration of external knowledge sources into the reasoning process. To address this limitation, recent research has proposed incorporating diverse knowledge sources into reasoning pipelines Tan et al. (2025a); Sun et al. (2023); Sarmah et al. (2024); Tan et al. (2025b); Su et al. (2023); Ma et al. (2023).
@@ 3.4 Retrieval Module
The retrieval module is responsible for extracting relevant information based on the given instruction and parent nodes. To eliminate training costs and enhance generalizability, we leverage off-the-shelf models to handle multimodality. However, retrieval across different source types can exhibit inherently different similarity score distributions, as we show in detail in Appendix A.1. To address this, we propose tailored strategies for different modalities. We firstly utilizes LLM to decompose the instruction into text-related and image-related components:
@@ 4.2 Main Results
To assess our method in long-range reasoning, we compare its performance with an iterative framework, SKURG in Figure 4, which shows the F1 score across varying numbers of required supporting contexts in MultimodalQA. As the number of supporting contexts increases, requiring more steps, SKURG’s performance drops significantly, especially beyond five contexts. In contrast, our model maintains stable performance, which shows its robustness in complex reasoning scenarios.
@@ 4.3 Ablation Study
We present the ablation studies for Adaptive Planning Graph-guided planning , triplet conversion and retrieval module on MultimodalQA. For single-hop tasks, removing graph-guided mechanism results in minor performance decline. However, for multi-hop questions, it leads to a substantial drop of over 10 points in both F1 and exact match scores. It demonstrates that our planning module has a critical role in handling questions that require long-range inference steps. Eliminating triplet conversion during knowledge base construction results in consistent drops across all metrics, demonstrating its role in aligning data granularity. For the retrieval module, we observe a more significant performance drop without our modality-specific strategies, as incorrect retrieval intensifies hallucination for both reasoning and global planning. It validates that our retrieval module improve the performance even without additional training.
@@ 4.4 Case Study
We show a case study in Figure 5, which presents how the Adaptive Planning Graph is constructed. At the initial few steps, the system attempted to retrieve information about Television of Tyler Alvarez. However, it incorrectly shifted the focus to the TV show Veronica Mars. Upon recognizing that this show did not align with the poster description in the question, the system redirected its reasoning toward Bailee Madison and relevant TV shows about the poster. The case study highlights that, even when the graph temporarily deviates, the system is capable of recovering and producing accurate results.
@@ 6 Limitations
7 Acknowledgment
This research was supported by the CSIRO Tech4HSE project RG242446. Xiaoyang Wang is supported by the Australian Research Council DP230101445 and DP240101322.
@@ Appendix A Appendices
A.1 Challenges of Modality-Agnostic Retrieval
Figure 6: Comparison of similarity score distributions.
Figure 7: Projection of image and text embeddings.
A.2 Overall Workflow
A.3 Detailed Prompts
@@ A.5 Implementation
In our training-free framework, models can be easily incorporated as required, without the need for fine-tuning. Here, we utilize Llama3-70B-Instruct as the LLM and LLaVA-13B as the visual-language model. As outlined in Section 3.2, the text knowledge base supports search within a text embedding space, while the image knowledge base expects cross-modal retrieval between text queries and images. To facilitate these retrieval processes, we employ CLIP to generate embeddings for both texts and images. During knowledge base construction, we utilize sklearn.BallTree structure to efficiently store and organize the embeddings. To retrieve candidates, we employ query_\_radius method, which returns all neighbors within a predefined distance threshold. We choose this method over top-$k$ retrieval as top-$k$ assumes a fixed number of relevant items, which is unsuitable for multimodal multi-hop question answering tasks, where the number of relevant sources can vary depending on the question.